# Frustratingly Simple Retrieval Improves Challenging, Reasoning-Intensive Benchmarks

### v1

@@ Abstract
Retrieval-augmented Generation (RAG) has primarily been studied in limited settings, such as factoid question answering; more challenging, reasoning-intensive benchmarks have seen limited success from minimal RAG. In this work, we challenge this prevailing view on established, reasoning-intensive benchmarks: MMLU, MMLU Pro, AGI Eval, GPQA, and MATH. We identify a key missing component in prior work: a usable, web-scale datastore aligned with the breadth of pretraining data. To this end, we introduce CompactDS: a diverse, high-quality, web-scale datastore that achieves high retrieval accuracy and subsecond latency on a single-node. The key insights are (1) most web content can be filtered out without sacrificing coverage, and a compact, high-quality subset is sufficient; and (2) combining in-memory approximate nearest neighbor (ANN) retrieval and on-disk exact search balances speed and recall. Using CompactDS, we show that a minimal RAG pipeline achieves consistent accuracy improvements across all benchmarks and model sizes (8B–70B), with relative gains of 10% on MMLU, 33% on MMLU Pro, 14% on GPQA, and 19% on MATH. No single data source suffices alone, highlighting the importance of diversity of sources (web crawls, curated math, academic papers, textbooks). Finally, we show that our carefully designed in-house datastore matches or outperforms web search engines such as Google Search, as well as recently proposed, complex agent-based RAG systems—all while maintaining simplicity, reproducibility, and self-containment. We release CompactDS and our retrieval pipeline, supporting future research exploring retrieval-based AI systems.
@@ 1 Introduction
To this end, we introduce CompactDS, a 380-billion-word datastore constructed from a high-quality, diverse collection of sources—including web crawls, curated math content, academic papers, and textbooks—designed to match the breadth of pretraining data while remaining accessible. CompactDS is built on two key insights. First, it is possible to aggressively filter low-quality web text while preserving the coverage and diversity of the web, resulting in a smaller yet representative dataset. Second, combining in-memory approximate nearest neighbor (ANN) search (johnson2019billion, ) with on-disk exact inner product search subramanya2019diskann enables subsecond retrieval on a single 456GB RAM node—far more efficient than prior systems that require several TBs of RAM.
CompactDS leads to significant performance improvements on benchmarks that were previously thought not to benefit from mininal retrieval: +10% on MMLU, +33% on MMLU Pro, +14% on GPQA, and +19% on MATH with LLaMa 3.1 8B Instruct, with persistent gains as the model scales to 70B parameters. Notably, exact inner product search with a more expressive encoder following ANN helps compared to using ANN alone, widening gains over no retrieval from 26% to 33% on MMLU Pro and from 14% to 19% on MATH. Our ablation study on datastore composition reveals several key insights: (1) Dataset diversity is critical: no single data source suffices alone, and even removing weakest contributing sources leads to degraded performance; (2) Including educational materials such as textbooks, often absent from web crawls, consistently improves performance; and (3) Quality filtering methods developed for pretraining generally help.
Notably, our in-house datastore matches or outperforms commercial search engines like Google Search across all benchmarks—a result that was previously unattainable using standard RAG benchmarks built with commercial search as an oracle (kwiatkowski2019natural, ). It also matches or outperforms recently proposed, complex systems based on commercial search and agentic pipelines (li2025search, ), as demonstrated on GPQA Diamond (rein2023gpqagraduatelevelgoogleproofqa, ) and MATH-500 (hendrycks2021measuringmathematicalproblemsolving, ) using QwQ 32B.
@@ 3.1 CompactDS Data Sources
We then augment the web crawl with data sources commonly regarded as high quality for pretraining, including Wikipedia and books (devlin2019bert, ), as well as textbooks (shi2025flexolmo, ).
@@
  • Textbooks: We take the data from shi2025flexolmo , text extracted from digitized PDF textbooks.
@@ 4.2 Main Results
Table 1 also reports the performance of datastores constructed from individual data sources. We find that certain sources greatly benefit specific benchmarks, e.g., Textbooks improve MMLU and GPQA; Math corpora boost performance on MATH; Wikipedia from DPR helps GPQA Biology; and Pes2o improves GPQA Chemistry. However, gains from any single source are generally limited.
@@
Among single-source datastores, educational content like textbooks, which is often absent in web-crawled corpora, and expert content like Math deliver the greatest improvement. It is also worth noting that Wikipedia from DPR is the most commonly used datastore in the RAG literature karpukhin-etal-2020-dense ; lewis2020retrieval however here, it offers little benefit on average, and even hurts performance on many datasets.
@@ 5.2 Results
PDFs, in particular, are a notable advantage of web search. While CompactDS includes PDFs, they are primarily academic papers and textbooks. In contrast, the web contains many valuable PDFs, such as lecture notes and problem set solutions, that are largely absent from CompactDS. To assess the impact of different source types, we compare performance using only web pages vs. only PDFs (Table 7). PDF-only retrieval can significantly improve performance on some tasks, e.g., on MMLU Pro, using only PDFs performs nearly as well as using both PDFs and web pages, and for GPQA, PDFs alone even outperform the combined. Overall, we think:
@@
    PDFs are a valuable source of information complementary to web crawl, similar to how textbooks and academic papers were effective in §4.2.
@@ C.2 Data source Leave-one-out
To further investigate the impact of individual data sources from §4.2, we conduct leave-one-out ablations where the retrieved passages from one data source are excluded at a time. Table 13 reports the complete results. For instance, MATH drops 2.6% when Math is excluded, MMLU Pro drops 1.6% when Textbooks is excluded, and GPQA Bio drops 3.9% when PubMed is excluded. This implies that the diversity of the data source is crucial. However, the performance can sometimes improve when excluding data sources, e.g., GPQA Bio improves by 5.1% after excluding High-quality CC, indicating that the retrieval can sometimes pick up noisy passages.

### v2

@@ Abstract
Retrieval-augmented Generation (RAG) has primarily been studied in limited settings, such as factoid question answering; more challenging, reasoning-intensive benchmarks have seen limited success from minimal RAG. In this work, we challenge this prevailing view on established, reasoning-intensive benchmarks: MMLU, MMLU Pro, AGI Eval, GPQA, and MATH. We identify a key missing component in prior work: a usable, web-scale datastore aligned with the breadth of pretraining data. To this end, we introduce CompactDS: a diverse, high-quality, web-scale datastore that achieves high retrieval accuracy and subsecond latency on a single-node. The key insights are (1) most web content can be filtered out without sacrificing coverage, and a compact, high-quality subset is sufficient; and (2) combining in-memory approximate nearest neighbor (ANN) retrieval and on-disk exact search balances speed and recall. Using CompactDS, we show that a minimal RAG pipeline achieves consistent accuracy improvements across all benchmarks and model sizes (8B–70B), with relative gains of 10% on MMLU, 33% on MMLU Pro, 14% on GPQA, and 19% on MATH. No single data source suffices alone, highlighting the importance of diversity of sources (web crawls, curated math, academic papers, educational text). Finally, we show that our carefully designed in-house datastore matches or outperforms web search engines such as Google Search, as well as recently proposed, complex agent-based RAG systems—all while maintaining simplicity, reproducibility, and self-containment. We release CompactDS and our retrieval pipeline, supporting future research exploring retrieval-based AI systems.
@@ 1 Introduction
To this end, we introduce CompactDS, a 380-billion-word datastore constructed from a high-quality, diverse collection of sources—including web crawls, curated math content, academic papers, and educational text—designed to match the breadth of pretraining data while remaining accessible. CompactDS is built on two key insights. First, it is possible to aggressively filter low-quality web text while preserving the coverage and diversity of the web, resulting in a smaller yet representative dataset. Second, combining in-memory approximate nearest neighbor (ANN) search (johnson2019billion, ) with on-disk exact inner product search subramanya2019diskann enables subsecond retrieval on a single 456GB RAM node—far more efficient than prior systems that require several TBs of RAM.
CompactDS leads to significant performance improvements on benchmarks that were previously thought not to benefit from mininal retrieval: +10% on MMLU, +33% on MMLU Pro, +14% on GPQA, and +19% on MATH with LLaMa 3.1 8B Instruct, with persistent gains as the model scales to 70B parameters. Notably, exact inner product search with a more expressive encoder following ANN helps compared to using ANN alone, widening gains over no retrieval from 26% to 33% on MMLU Pro and from 14% to 19% on MATH. Our ablation study on datastore composition reveals several key insights: (1) Dataset diversity is critical: no single data source suffices alone, and even removing weakest contributing sources leads to degraded performance; (2) Including educational materials, often absent from web crawls, consistently improves performance; and (3) Quality filtering methods developed for pretraining generally help.
Notably, our in-house datastore matches or outperforms web search engines like Google Search across all benchmarks—a result that was previously unattainable using standard RAG benchmarks built with web search as an oracle (kwiatkowski2019natural, ). It also matches or outperforms recently proposed, complex systems based on web search and agentic pipelines (li2025search, ), as demonstrated on GPQA Diamond (rein2023gpqagraduatelevelgoogleproofqa, ) and MATH-500 (hendrycks2021measuringmathematicalproblemsolving, ) using QwQ 32B.
@@ 3.1 CompactDS Data Sources
We then augment the web crawl with data sources commonly regarded as high quality for pretraining, including Wikipedia and books (devlin2019bert, ), as well as educational text (shi2025flexolmo, ).
@@
  • Educational Text: We take the data from shi2025flexolmo , educational text extracted from digitized PDFs.
@@ 4.2 Main Results
Table 1 also reports the performance of datastores constructed from individual data sources. We find that certain sources greatly benefit specific benchmarks, e.g., Educational Text improves MMLU and GPQA; Math corpora boost performance on MATH; Wikipedia from DPR helps GPQA Biology; and Pes2o improves GPQA Chemistry. However, gains from any single source are generally limited.
@@
Among single-source datastores, educational content, which is often absent in web-crawled corpora, and expert content like Math deliver the greatest improvement. It is also worth noting that Wikipedia from DPR is the most commonly used datastore in the RAG literature karpukhin-etal-2020-dense ; lewis2020retrieval however here, it offers little benefit on average, and even hurts performance on many datasets.
@@ 5.2 Results
PDFs, in particular, are a notable advantage of web search. While CompactDS includes PDFs, they are primarily academic papers and educational text. In contrast, the web contains many valuable PDFs, such as lecture notes and problem set solutions, that are largely absent from CompactDS. To assess the impact of different source types, we compare performance using only web pages vs. only PDFs (Table 7). PDF-only retrieval can significantly improve performance on some tasks, e.g., on MMLU Pro, using only PDFs performs nearly as well as using both PDFs and web pages, and for GPQA, PDFs alone even outperform the combined. Overall, we think:
@@
    PDFs are a valuable source of information complementary to web crawl, similar to how educational text and academic papers were effective in §4.2.
@@ C.2 Data source Leave-one-out
To further investigate the impact of individual data sources from §4.2, we conduct leave-one-out ablations where the retrieved passages from one data source are excluded at a time. Table 13 reports the complete results. For instance, MATH drops 2.6% when Math is excluded, MMLU Pro drops 1.6% when Educational Text is excluded, and GPQA Bio drops 3.9% when PubMed is excluded. This implies that the diversity of the data source is crucial. However, the performance can sometimes improve when excluding data sources, e.g., GPQA Bio improves by 5.1% after excluding High-quality CC, indicating that the retrieval can sometimes pick up noisy passages.